# Evaluating the Generalization Ability of Spatiotemporal Model in Urban Scenario

### v1

@@ 1 Introduction
2 Related Work
2.1 Spatiotemporal Benchmarks
@@ 3 Problem Formulation
4 The ST-OOD Benchmark
This section formally introduces the proposed ST-OOD benchmark, designed to comprehensively evaluate the spatiotemporal generalization capabilities of current models across various scenarios. We begin by detailing the collection and organization of ST-OOD in Sec. 4.1. The detailed information is listed on Table I and the node distribution is visualized in Figure 2. Then, in Sec. 4.2, we conduct an in-depth data analysis to gain deeper insights into ST-OOD.
@@ 4 The ST-OOD Benchmark
4.1 Data Collection and Preprocessing
@@ 5 Experiments
D2STGNN 57% 55% 45% 57% 11% 39% 53% 106% 24% 25% 36% 40% 3.83 5.17
STID 60% 56% 53% 57% 9% 25% 64% 116% 26% 24% 39% 40% 6.33 5.17
MLP 63% 62% 51% 56% 10% 9% 73% 83% 27% 18% 52% 54% 11.50 5.33
LSTM 60% 60% 48% 54% 11% 10% 66% 82% 26% 18% 49% 54% 8.83 5.33
WaveNet 62% 61% 50% 55% 11% 10% 66% 85% 27% 18% 44% 48% 10.00 5.33
STAEformer 58% 55% 51% 62% 9% 29% 66% 127% 25% 25% 34% 37% 4.00 6.00
GWNet 56% 54% 50% 61% 10% 21% 63% 129% 26% 26% 40% 44% 4.50 6.50
STGODE 60% 57% 48% 58% 10% 34% 66% 117% 26% 25% 36% 38% 6.67 6.50
STTN 59% 56% 48% 57% 10% 25% 96% 194% 27% 27% 35% 46% 6.83 7.83
MTGNN 58% 57% 46% 60% 10% 50% 66% 131% 26% 27% 35% 38% 4.33 8.83
STGCN 61% 57% 50% 59% 10% 81% 97% 235% 24% 27% 35% 38% 8.00 9.17
DAAGCN 59% 57% 48% 63% 10% 44% 80% 156% 26% 26% 43% 45% 7.33 9.33
AGCRN 60% 57% 48% 59% 10% 51% 109% 255% 27% 26% 42% 47% 8.83 10.50
@@
STID 60% 56% 53% 57% 9% 25% 64% 116% 26% 24% 39% 40%
w.o. Node 63% 62% 51% 56% 10% 9% 73% 83% 27% 18% 52% 54%
CaST 86% 64% 68% 73% 23% 43% 88% 96% 35% 114% 74% 98%
CauSTG 95% 71% 75% 80% 25% 47% 96% 105% 38% 125% 82% 108%
STONE 65% 51% 51% 58% 17% 34% 66% 76% 27% 90% 56% 78%
@@ 6 Discussion
7 Future Opportunities and Limitations

### v2

@@ I Introduction
II Related Work
II-A Spatiotemporal Benchmarks
@@ III Problem Formulation
IV The ST-OOD Benchmark
This section formally introduces the proposed ST-OOD benchmark, designed to comprehensively evaluate the spatiotemporal generalization capabilities of current models across various scenarios. We begin by detailing the collection and organization of ST-OOD in Sec. IV-A. The detailed information is listed on Table I and the node distribution is visualized in Figure 2. Then, in Sec. IV-B, we conduct an in-depth data analysis to gain deeper insights into ST-OOD.
@@ IV The ST-OOD Benchmark
IV-A Data Collection and Preprocessing
@@ V Experiments
D2STGNN 57.41% 54.87% 44.88% 56.80% 10.58% 38.75% 52.99% 105.76% 23.76% 25.42% 35.74% 39.96% 3.83 5.17
STID 60.14% 56.36% 53.31% 56.85% 9.34% 25.49% 64.38% 115.60% 25.87% 24.40% 38.74% 40.06% 6.33 5.17
MLP 63.27% 61.89% 51.02% 55.79% 10.33% 9.24% 72.83% 82.55% 27.21% 18.34% 51.75% 53.72% 11.50 5.33
LSTM 60.48% 59.50% 47.94% 54.01% 10.57% 9.56% 65.80% 82.45% 26.45% 18.50% 49.37% 54.05% 8.83 5.33
WAVENET 61.60% 60.84% 49.55% 54.92% 10.56% 9.50% 66.46% 84.69% 26.74% 18.40% 43.97% 47.93% 10.00 5.33
STAEFORMER 57.58% 55.37% 50.70% 61.87% 8.86% 28.65% 65.60% 127.35% 24.96% 24.84% 33.81% 36.63% 4.00 6.00
GWNET 55.56% 53.95% 49.86% 61.25% 9.78% 21.08% 63.04% 128.80% 25.54% 26.11% 40.11% 43.56% 4.50 6.50
STGODE 60.40% 57.20% 48.40% 58.02% 10.13% 34.44% 65.54% 117.00% 26.38% 24.65% 35.78% 37.68% 6.67 6.50
STTN 59.07% 56.25% 48.27% 57.41% 9.84% 25.42% 96.46% 194.48% 27.45% 27.30% 35.23% 45.88% 6.83 7.83
MTGNN 58.45% 56.69% 45.78% 59.77% 10.05% 49.68% 65.76% 130.81% 25.93% 27.34% 35.09% 37.56% 4.33 8.83
STGCN 61.36% 56.60% 50.16% 58.99% 10.30% 81.22% 96.83% 234.73% 24.20% 27.19% 35.33% 38.01% 8.00 9.17
DAAGCN 58.55% 56.51% 48.43% 63.17% 10.04% 43.92% 79.84% 156.02% 26.26% 26.41% 43.38% 44.63% 7.33 9.33
AGCRN 59.98% 57.12% 48.39% 59.10% 10.16% 51.36% 108.95% 254.95% 26.89% 26.49% 41.72% 47.32% 8.83 10.50
@@
STID 60.14% 56.36% 53.31% 56.85% 9.34% 25.49% 64.38% 115.60% 25.87% 24.40% 38.74% 40.06%
w.o. Node 63.27% 61.89% 51.02% 55.79% 10.33% 9.24% 72.83% 82.55% 27.21% 18.34% 51.75% 53.72%
CaST 86.24% 64.73% 68.91% 73.56% 23.18% 43.05% 88.39% 96.82% 35.47% 114.69% 74.31% 98.95%
CauSTG 95.63% 71.09% 75.42% 80.17% 25.84% 47.36% 96.51% 105.28% 38.75% 125.93% 82.04% 108.70%
STONE 65.12% 51.79% 51.38% 58.96% 17.25% 34.61% 66.87% 76.43% 27.59% 90.21% 56.34% 78.08%
@@ VI Discussion
VII Future Opportunities and Limitations